OpenStudy (anonymous):

find the equation of a circle that has it's center at (-4,3) and has a radius of 5

OpenStudy (btaylor):

general form: (x-h)^2+(y-k)^2 = r^2 [the center is (h,k)] so: (x+4)^2 + (y-3)^2 = 25

OpenStudy (anonymous):

thanks